The firms production function is given as f(L,K) = 4L0.5K0.5 and input prices are w = 4 and r = 16. In the short-run, the level of capital is fixed at K = 2. What is the firms short-run average cost function? Wählen Sie eine Antwort: O a. 0.125q 32/q b. 0.25q 16/q О с. 0.125q O d, 4 + 32/q
Answer #1

Given, production function  0.5 -0.5

Cost of Labour, w = 4

Cost of capital, r = 16

Short run capital is fixed at K = 2

First of all determine the marginal product of Labour. Differentiate Q wrt L we get

dL dL

dL

K is independent of L so the derivative K (wrt L) = 0

large MP_L =4*0.5K^{0.5}L^{0.5-1}

0.5 T 0.5 MPL = 0.5 * AKosLos

M PL 0.5 *

Similarly differentiating Q wrt K we get the marginal product of capital

MPK 0.5

Now according to equimarginal criterion satisfy the following formula

MPL w

0.5Q 0.5 16 16

large rac{K}{L} =rac{1}{4}

Therefore, K = 0.25L

Given, K = 2

0.5 -0.5

large Q = 4(L*0.25L)^{0.5}

Q = 4025L2)0.5

large Q = 4*0.5*L

large Q = 2L

Therefore, 4L = 2Q

The cost function can be written as

large C = rK + wL

C = 16K + 4L

Average cost function

AC

large AC = rac{16*2 + 4L}{Q}

large AC = rac{32}{Q} + rac{2Q}{Q}

large AC = 2 + rac{32}{Q}

So it is AC = 2 + 32/Q

I have tried all the methods but I am getting this answer only.